OpenStudy (anonymous):

Which of the following recursive formulas represents the same arithmetic sequence as the explicit formula an = 2 + (n - 1)5?

OpenStudy (anonymous):

A. an = (an - 1 + 5) 2 B. an = an - 1 5 C. an = an - 1 + 2 D. an = an - 1 + 5

OpenStudy (anonymous):

@SolomonZelman